OpenStudy (geometry_hater):

Multiply this matrix [2 1 2] [-1 0] [3 5 4] X [2 4] [1 0 3] [3 3]

OpenStudy (mathmate):

The element c11 is obtained by summing the three products a11 b11 + a12 b21 + a13 b31 =2(-1) + 1(2) + 2(3) =-2+2+6 =6 Continue with the remaining 5 elements.

OpenStudy (geometry_hater):

uhh where did the a 11 b 11 come from? I thought your suppose to multiply the collum times the rom BTW this is Algebra II

OpenStudy (geometry_hater):

*I meant column X row

OpenStudy (mathmate):